The pediatric infusion rate for potassium is 5 mEq/h. If 9 mL of a 39.2% solution of potassium acetate (KC2H3O2) is diluted to 1 L of infusion solution, calculate the proper infusion rate in mL/h. The correct answer is 138.89 mL/h potassium acetate infusion. Please show how this question is solved.
Answer #1

Molecular weight= Atomic mass of the element *No. of atoms of that element

( add up all the molecular weight of the elements )

Molecular weight of potassium acetete(KC2H3O2)=1*39.0983+2*12.0107+3*1.00794+2*15.9994=98.1423 g/mol

Atomic mass of

  • Carbon: 12.0107
  • Potassium: 39.0983
  • Hydrogen: 1.00794
  • Oxygen : 15.9994

Concentration of potassium acetate* volume given = Concentration of potassium acetate after dilution *1000

39.2*9=C*1000

C=352.8/1000

Concentration of potassium actetate diluted infusion solution =0.353%

Concentration of potassium actetate diluted infusion solution in 100 ml =0.353g/100ml=3.53mg/ml

The infusion rate of diluted solution with potassium infused at 5 mEq/h = (3.53*98.1423*2)/5=692.884638/5=138.58 ml/hr

(valence of potassium acetate = 2)